OpenStudy (anonymous):

can some one please help me out with and example of a subtraction operation on function.

OpenStudy (anonymous):

\[f(x)=e^x\] \[g(x)=\sqrt{x}\] \[(f-g)(x)=e^x-\sqrt{x}\]